# Super-LIO: A Robust and Efficient LiDAR-Inertial Odometry System with a Compact Mapping Strategy

### v1

@@ III-C Octo-Voxel-Based Map (OctVox)
Data structure. We employ a hashed voxel grid with a least-recently-used (LRU) policy for memory management. The map is defined as $\mathcal{M}=\{\mathcal{V}_{i}\mid i\in\mathbb{N}\}$, that is, a collection of voxels. Each voxel $\mathcal{V}_{i}$ of edge length $r_{v}$ is subdivided into $2\times 2\times 2$ subvoxels $\mathcal{V}_{i,s}$ of edge length $r_{\text{s}}=\tfrac{1}{2}r_{v}$. A voxel allocates eight contiguous entries, each storing a representative $\bm{\mu}_{i,s}\in\mathbb{R}^{3}$ and a counter $n_{i,s}$. The structure can be expressed as
$\mathcal{V}_{i}=\{\mathcal{V}_{i,s}\mid s\in[0,7]\},\ \ \mathcal{V}_{i,s}=(\bm{\mu}_{i,s},n_{i,s}).$
This contiguous layout enforces an explicit density cap of eight representatives per voxel and achieves noise suppression in the global frame through incremental averaging within each subvoxel. The hash table uses Robin Hood hashing with open addressing [32] and is implemented with tsl::robin_map [33], improving cache locality and providing near-constant lookup time. When the memory budget is reached, the LRU policy evicts the least-recently-accessed voxels.
Voxel indexing and update. Given a point $\hat{\mathbf{p}}^{\mathcal{G}}\in\mathbb{R}^{3}$ in the world frame, we quantize it to subvoxel resolution using Eq.(4) to obtain its parent voxel key $\mathbf{k}\in\mathbb{Z}^{3}$ and the local linear subvoxel index $s\in[0,7]$.
$\mathbf{k}^{\text{sub}}=\Big{\lfloor}\tfrac{\hat{\mathbf{p}}^{\mathcal{G}}}{r_{\text{s}}}\Big{\rfloor},\quad\mathbf{k}=\mathbf{k}^{\text{sub}}\mathbin{\text{{>>}}}1,$ (4)
@@
Here, &\& and $|$ denote bitwise AND/OR, while $\mathbin{\text{{<<}}}$ and $\mathbin{\text{{>>}}}$ denote left and right shifts, respectively. Equations (4) therefore provide both the voxel key $\mathbf{k}$ and the subvoxel index $s$ in $O(1)$ time, using bitwise operations without conditional branching.
Once $(\mathbf{k},s)$ is determined, the corresponding record $(\bm{\mu}_{s},n_{s})$ is updated incrementally. If the entry is uninitialized, it is set to $(\hat{\mathbf{p}}^{\mathcal{G}},1)$. Otherwise, if $\left\lVert\hat{\mathbf{p}}^{\mathcal{G}}-\bm{\mu}_{s}\right\rVert_{2}\leq\tau_{\text{merge}}$ and $n_{s}\leq n_{\text{max}}$, the update rule
$\bm{\mu}_{s}\leftarrow\bm{\mu}_{s}+\tfrac{1}{n_{s}+1}\big{(}\hat{\mathbf{p}}^{\mathcal{G}}-\bm{\mu}_{s}\big{)},\quad n_{s}\leftarrow n_{s}+1,$ (5)
is applied. This corresponds to an unbiased incremental mean update, with the estimation variance asymptotically decreasing as $1/n_{s}$.
@@ III-D Heuristic-Guided KNN Search
A major bottleneck in LiDAR-inertial odometry is the repeated KNN search for scan-to-map residuals. Even with OctVox map that cap per-voxel density, a naive search must examine all representatives within a ball of radius $R$. Small $R$ often yields unstable neighbors in sparse areas, whereas larger $R$ improves quality but requires visiting $\Theta((R/r_{v})^{3})$ voxels, so the candidate set grows cubically with $R$. Tree-based indices (e.g., KD-trees) scale as $\mathcal{O}(\log N)$ with map size, but lack the constant-time locality of voxel hashing. To address this, we design an HKNN that groups candidate subvoxels by distance and exploits octant symmetry to unify traversal, enabling efficient and robust correspondence search.
Figure 3: Illustration of the heuristic KNN candidate selection process based on subvoxel proximity. A query subvoxel (blue) is surrounded by multiple neighboring subvoxels at increasing distances, which are color-coded by their Manhattan distance to the center. These candidates are incrementally inserted into a priority queue and sorted based on their geometric proximity to the query, allowing for an efficient Top-K retrieval.
HKNN list generation. As shown in Fig.3, the HKNN list is defined as an ordered sequence of subvoxel groups, each group contains candidate subvoxels clustered by their spatial distance from the query subvoxel. The list is generated once during initialization from the maximum search radius and voxel resolution and remains fixed throughout operation. In contrast to conventional neighborhood definitions[15, 18, 17] that require exhaustive traversal of a fixed set, the HKNN list enables group-wise traversal with early termination.
@@
Let $\mathbf{K}=\{\mathbf{k}_{i}\}$ denote the set of voxel indices within the maximum search radius $R_{\max}$, where each index $\mathbf{k}_{i}$ is defined relative to the query voxel center at the origin, $\mathbf{k}_{0}=(0,0,0)$. The corresponding set of subvoxels is given by $\mathbf{S}\;=\;\mathbf{K}\times\{0,\ldots,7\}$. Accordingly, each element can be expressed as $\mathbf{S}_{i,j}=(\mathbf{k}_{i},s_{j})$.
Due to the discrete nature of the voxel grid, the distances between the reference subvoxel $\mathbf{S}_{0,0}$ and all subvoxels $\mathbf{S}_{i,j}\in\mathbf{S}$, computed by Eq. (6), constitute a finite enumerated set of attainable distances. We denote this set by $\mathcal{D}=\{d_{m}\mid m=0,\dots,M\}$, which is arranged in ascending order. Subvoxels at the same distance are then grouped into a distance-equivalence set, as defined in Eq.(7), where $\mathcal{V}(\mathbf{S}_{i,j})$ denotes the cubic region corresponding to subvoxel $\mathbf{S}_{i,j}$.
$\mathcal{H}_{m}\;=\;\big{\{}\,\mathbf{S}_{i,j}\in\mathbf{S}\;:\;\mathrm{dis}\big{(}\mathcal{V}(\mathbf{S}_{0,0}),\mathcal{V}(\mathbf{S}_{i,j})\big{)}=d_{m}\,\big{\}}$ (7)
Concatenating the groups yields the ordered traversal list defined in Eq. (8).
@@
Although in the absolute coordinate frame each subvoxel $\mathbf{S}_{0,j}$ appears to have a distinct traversal list, in the relative coordinate frame all subvoxels share the same search topology. They are related by reflections along the three coordinate axes, a property we term octant symmetry. This symmetry implies that the traversal list for any $\mathbf{S}_{0,j}$ can be obtained from the reference list $\mathcal{T}(\mathbf{S}_{0,0})$ by a deterministic transformation.
According to the subvoxel indexing rule in Eq. (4), for each entry in the reference traversal list Eq. (8) and for any local subvoxel, the corresponding voxel index and subvoxel index are computed as Eq. (9), where $(k_{x},k_{y},k_{z},s)$ denotes the entry in $\mathcal{T}(\mathbf{S}_{0,0})$, $s_{j}$ is the query subvoxel index, $(b_{x},b_{y},b_{z})$ are the intermediate variables of $s_{j}$ defined in Eq. (4), and $\oplus$ denotes the bitwise XOR.
$\mathbf{k}^{\prime}=\big{(}(-1)^{b_{x}}k_{x},\;(-1)^{b_{y}}k_{y},\;(-1)^{b_{z}}k_{z}\big{)},\ \ s^{\prime}=s\oplus s_{j}.$ (9)
@@
The detailed procedure is summarized in Algorithm 1, where $|\mathbf{H}|$ and $\mathrm{maxdist}(\mathbf{H})$ denote the number of elements in the max-heap $\mathbf{H}$ and the largest distance among them, respectively.
Input: Query point $\mathbf{p}^{\mathcal{I}}$ (IMU frame); prior state $\hat{X}^{-}$; voxel map $\mathcal{M}$; canonical traversal list $\mathcal{T}(\mathbf{S}_{0,0})$ Per Eq. (8); group lower bounds $\mathcal{D}=\{d_{m}\}_{m=0}^{M}$; fixed radius $R\leq R_{max}$; target $K$
@@
2Transform $\mathbf{p}^{\mathcal{I}}$ to $\hat{\mathbf{p}}^{\mathcal{G}}$ using $\hat{X}^{-}$ (Eq. (3)).
@@
8 if $|\mathbf{H}|=K$ and $d_{m}>\mathrm{maxdist}(\mathbf{H})$ then break;
@@
13    $\Delta\mathbf{k}\leftarrow(\sigma_{x}k_{x},\ \sigma_{y}k_{y},\ \sigma_{z}k_{z})$
14    $\mathbf{k}^{\prime}\leftarrow\mathbf{k}+\Delta\mathbf{k}$  and  $s^{\prime}\leftarrow s\oplus s_{p}$
15    $\bm{\mu}\leftarrow\mathcal{V}(\mathbf{k}^{\prime},s^{\prime})$
16    $\delta\leftarrow\big{\|}\hat{\mathbf{p}}^{\mathcal{G}}-\bm{\mu}\big{\|}_{2}$
@@ 15    $\bm{\mu}\leftarrow\mathcal{V}(\mathbf{k}^{\prime},s^{\prime})$
III-E Discussion
We use PCA on neighboring points to fit local planes and form point-to-plane residuals, a formulation introduced in [7] and widely validated in subsequent LIO systems for stable performance. Our proposed map structure and the corresponding HKNN search are highly efficient for LiDAR SLAM and can be easily integrated into existing systems.
Compared with voxel-neighbor methods [15], enlarging the association range usually comes with high cost. Our HKNN offers better tolerance to the choice of $R$ and yields more reliable neighbors in sparse scans, but excessively increasing $R$ still introduces unnecessary overhead, especially in LiDAR-inertial odometry applications.
Empirically, the hyperparameter $\tau_{\text{merge}}$ (Section III-C) is most effective for non-repetitive LiDARs, as it constrains the averaging range and mitigates single-shot noise, while its impact on mechanical spinning LiDARs is minor. Setting it to about $3\sigma$ of the noise is generally reasonable. We note this to guide users of our open-source implementation in parameter tuning; for all experiments in Section IV, we used the same configuration across sensors and environments.
@@ IV EXPERIMENT
In this section, we evaluate the proposed Super-LIO system on two computing platforms, X86 and ARM, as well as on multiple datasets. We compare its performance against state-of-the-art lightweight LIO frameworks. The evaluation covers both accuracy and efficiency, with a particular emphasis on runtime performance under resource-constrained conditions. We further conduct robustness tests across a wide variety of real-world scenarios.
@@ IV-A Experimental Setup
We compare Super-LIO against several lightweight LiDAR-Inertial Odometry baselines, including FAST-LIO2 [14], Faster-LIO [15], and iG-LIO [18]. To isolate the contribution of our HKNN strategy, we further implement an ablation variant, Super-LIO*, which is identical to Super-LIO but replaces HKNN with the 18-neighbor voxel search of Faster-LIO. In addition, we provide a concurrency‑optimized version of FAST‑LIO2, denoted as FAST‑LIO2*, for fair real-time evaluation.
For fairness, all methods use identical parameter settings: maximum iterations = 4, random downsampling rate = 3, and voxel downsampling resolution = 0.5 m. For our method and other voxel-based approaches [15, 18], the map voxel size is also set to 0.5 m. The HKNN in Super-LIO is configured with $R_{\text{max}}=0.875$ m, covering a $7\times 7\times 7$ subvoxel region that offers a larger search range than Super-LIO* and Faster-LIO. Except for LiDAR–IMU extrinsic calibrations, which vary across datasets, all other parameters are kept at default values throughout the experiments.
Experiments are conducted on both public and private datasets. Public benchmarks include M2DGR [34], NCLT [35], MCD [36], and NTU [37], all with ground-truth trajectories for accuracy evaluation. Our private dataset comprises ten sequences covering diverse indoor and outdoor environments, used for runtime and robustness tests. All datasets are played at $5\times$ real-time speed on the X86 platform to assess both throughput and accuracy, and at $1\times$ speed on the Orin NX to evaluate real-time performance under resource constraints. Examples of private data collection platforms and sequences are shown in Fig. 4 and Fig. 5.
Figure 5: Ten self-collected point cloud sequences deliberately selected to cover diverse environment dimensions—indoor vs. outdoor, structured vs. unstructured, and open vs. confined—including parks and ponds, forests, drone-racing arenas, factories, underground garages, and office spaces. The sequences are indexed from 01 to 10 in row-major order (left to right, top to bottom), where 04, 05, 07, and 09 are collected by autonomous UAV flights.
TABLE I: RMSE (m) on Public Datasets (X86 @ 5× Speed)

### v2

@@ III-C Octo-Voxel-Based Map (OctVox)
Data structure. We employ a hashed voxel grid with a least-recently-used (LRU) policy for voxel caching [faster_lio]. The map is defined as $\mathcal{M}=\{\mathcal{V}_{i}\mid i\in\mathbb{N}\}$, i.e., a collection of voxels. Each voxel $\mathcal{V}_{i}$ of edge length $r_{v}$ is subdivided into $2\times 2\times 2$ subvoxels $\mathcal{V}_{i,s}$ of edge length $r_{s}=\tfrac{1}{2}r_{v}$. A voxel allocates eight contiguous entries, each storing a representative $\bm{\mu}_{i,s}\in\mathbb{R}^{3}$ and a counter $n_{i,s}$. The structure can be expressed as
$\mathcal{V}_{i}=\{\mathcal{V}_{i,s}\mid s\in\{0,\dots,7\}\},\ \ \mathcal{V}_{i,s}=(\bm{\mu}_{i,s},n_{i,s}).$
This contiguous layout enforces an explicit density cap of eight representatives per voxel and achieves noise suppression in the global frame through incremental averaging within each subvoxel. The hash table uses Robin Hood hashing with open addressing [celis1985robin] and is implemented with tsl::robin_map [tsl_robin_map], improving cache locality and providing near-constant lookup time. When the hash table reaches its maximum voxel capacity, the LRU policy evicts the least-recently-accessed voxels.
Voxel indexing and update. Given a point $\hat{\bm{p}}^{\mathcal{G}}\in\mathbb{R}^{3}$ in the world frame, we quantize it to subvoxel resolution using Eq.(4) to obtain its parent voxel key $\mathbf{k}\in\mathbb{Z}^{3}$ and the local linear subvoxel index $s$.
$\mathbf{k}^{\text{sub}}=\Big\lfloor\tfrac{\hat{\bm{p}}^{\mathcal{G}}}{r_{\text{s}}}\Big\rfloor,\quad\mathbf{k}=\mathbf{k}^{\text{sub}}\mathbin{\text{{>>}}}1,$ (4)
@@
Here, &\& and $|$ denote bitwise AND/OR, while $\mathbin{\text{{<<}}}$ and $\mathbin{\text{{>>}}}$ denote left and right shifts, respectively. Equations (4) therefore provide both the voxel key $\mathbf{k}$ and the subvoxel index $s$ in $O(1)$ time using bitwise operations.
After obtaining $(\mathbf{k},s)$, the corresponding subvoxel entry $(\bm{\mu}_{s},n_{s})$ is retrieved from the hash table. If uninitialized, it is set to $(\hat{\bm{p}}^{\mathcal{G}},1)$; otherwise, when $\lVert\hat{\bm{p}}^{\mathcal{G}}-\bm{\mu}_{s}\rVert_{2}\leq\tau_{\text{merge}}$ and $n_{s}\leq n_{\text{max}}$, an incremental mean update is applied as Eq. (5). This corresponds to an unbiased incremental mean update, with the estimation variance asymptotically decreasing as $1/n_{s}$.
$\bm{\mu}_{s}\leftarrow\bm{\mu}_{s}+\tfrac{1}{n_{s}+1}\big(\hat{\bm{p}}^{\mathcal{G}}-\bm{\mu}_{s}\big),\quad n_{s}\leftarrow n_{s}+1,$ (5)
@@ III-D Heuristic-Guided KNN Search
A major bottleneck in LiDAR-Inertial Odometry is the repeated KNN search for scan-to-map residuals. Even with OctVox map that cap per-voxel density, a naive search must examine all representatives within a ball of radius $R$. Small $R$ often yields unstable neighbors in sparse areas, whereas larger $R$ improves quality but requires visiting $\Theta((R/r_{v})^{3})$ voxels, so the candidate set grows cubically with $R$. Tree-based indices (e.g., KD-Tree) scale as $\mathcal{O}(\log N)$ with map size, but lack the constant-time locality of voxel hashing. To address this, we design an HKNN that groups candidate subvoxels by distance and exploits subvoxel symmetry to unify traversal, enabling efficient and robust correspondence search.
Figure 3: Illustration of the HKNN candidate selection process based on subvoxel proximity. A query subvoxel (blue) is surrounded by neighboring subvoxels at increasing distances, which are color-coded by the minimum Euclidean distance between their nearest vertices. These candidates are pre-grouped by geometric proximity; the groups are traversed sequentially in order of increasing distance until $K$ neighbors are found and no subsequent group can yield a closer neighbor, enabling efficient Top-$K$ retrieval without per-query sorting.
HKNN list generation. As shown in Fig.3, the HKNN list is defined as an ordered sequence of subvoxel groups, each group contains candidate subvoxels clustered by their spatial distance from the query subvoxel. The list is generated once during initialization from the maximum search radius and voxel resolution and remains fixed throughout operation. In contrast to conventional neighborhood definitions[faster_lio, ig_lio] that require exhaustive traversal of a fixed set, the HKNN list enables group-wise traversal with early termination.
@@
Let $\mathbf{K}=\{\mathbf{k}_{i}\}$ denote the set of voxel indices within the maximum search radius $R_{\max}$, where each index $\mathbf{k}_{i}$ is defined relative to the query voxel center at the origin, $\mathbf{k}_{0}=(0,0,0)$. The corresponding set of subvoxels is given by $\mathbf{S}\;=\;\mathbf{K}\times\{0,\ldots,7\}$, and each subvoxel element is indexed by $(\mathbf{k}_{i},s_{j})$ and denoted as $\mathbf{S}_{i,j}$.
Due to the discrete nature of the voxel grid, the distances between the reference subvoxel $\mathbf{S}_{0,0}$ and all subvoxels $\mathbf{S}_{i,j}\in\mathbf{S}$, computed by Eq. (6), constitute a finite enumerated set of attainable distances. We denote this set by $\mathcal{D}=\{d_{m}\mid m=0,\dots,M\}$, which is arranged in ascending order. Subvoxels at the same distance are then grouped into a distance-equivalence set, as defined in Eq.(7).
$\mathcal{H}_{m}\;=\;\big\{\,\mathbf{S}_{i,j}\in\mathbf{S}\;:\;\mathrm{dis}\big(\mathbf{S}_{0,0},\mathbf{S}_{i,j}\big)=d_{m}\,\big\}$ (7)
Concatenating the distance-equivalence groups yields the ordered traversal list for the reference subvoxel $\mathbf{S}_{0,0}$ of the query voxel, as defined in Eq. (8).
@@
Traversal lists for the remaining seven subvoxels are obtained by reflecting the reference list $\mathcal{T}(\mathbf{S}_{0,0})$ across the three coordinate axes. For each entry $(k_{i}^{x},k_{i}^{y},k_{i}^{z},s_{j})$ in $\mathcal{T}(\mathbf{S}_{0,0})$, the corresponding voxel and subvoxel indices $(\mathbf{k}^{\prime}_{i},s^{\prime}_{j})$ are computed as Eq. (9), where $s$ is the local subvoxel index, $(b_{x},b_{y},b_{z})$ are the bit variables of $s$ defined in Eq. (4), and $\oplus$ denotes bitwise XOR. According to Eq. (9), the symmetry operation leaves $\mathcal{T}(\mathbf{S}_{0,0})$ unchanged.
$\mathbf{k}^{\prime}_{i}=\big((-1)^{b_{x}}k_{i}^{x},\;(-1)^{b_{y}}k_{i}^{y},\;(-1)^{b_{z}}k_{i}^{z}\big),\quad s^{\prime}_{j}=s\oplus s_{j},$ (9)
@@
The detailed procedure is summarized in Algorithm 1, where $|\mathbf{H}|$ and $\mathrm{maxdist}(\mathbf{H})$ denote the number of elements in the max-heap $\mathbf{H}$ and the largest distance among them, respectively. $\mathcal{V}(\mathbf{k}_{g},s^{\prime})$ denotes the unique subvoxel identified by the global voxel index $\mathbf{k}_{g}$ and the subvoxel index $s^{\prime}$.
Input: Query point $\bm{p}^{\mathcal{I}}$ (IMU frame); prior state $\hat{\mathbf{X}}^{-}$; voxel map $\mathcal{M}$; canonical traversal list $\mathcal{T}(\mathbf{S}_{0,0})$ Per Eq. (8); group lower bounds $\mathcal{D}=\{d_{m}\}_{m=0}^{M}$; fixed radius $R\leq R_{max}$; target $K$
@@
2Transform $\bm{p}^{\mathcal{I}}$ to $\hat{\bm{p}}^{\mathcal{G}}$ using $\hat{\mathbf{X}}^{-}$ (Eq. (3)).
@@
8 if $|\mathbf{H}|=K$ and $d_{i}>\mathrm{maxdist}(\mathbf{H})$ then break;
@@
13    $\mathbf{k}^{\prime}\leftarrow(\sigma_{x}k_{x},\ \sigma_{y}k_{y},\ \sigma_{z}k_{z})$
14    $\mathbf{k}_{g}\leftarrow\mathbf{k}_{p}+\mathbf{k}^{\prime}$ and $s^{\prime}\leftarrow s\oplus s_{p}$
15    $\bm{\mu}\leftarrow\mathcal{V}(\mathbf{k}_{g},s^{\prime})\leftarrow(\mathbf{k}_{g},s^{\prime})$
16    $\delta\leftarrow\big\|\hat{\bm{p}}^{\mathcal{G}}-\bm{\mu}\big\|_{2}$
@@ IV Experiment
In this section, we evaluate Super-LIO on both X86 and ARM platforms and across multiple datasets, comparing it against state-of-the-art lightweight LIO frameworks. The experiments cover trajectory accuracy, per-platform runtime, module-level time analysis, and system resource usage such as CPU load and memory consumption.
@@ IV-A Experimental Setup
We compare Super-LIO against several lightweight LiDAR-Inertial Odometry baselines, including FAST-LIO2 [fast_lio2], Faster-LIO [faster_lio], and iG-LIO [ig_lio]. To isolate the contribution of the HKNN strategy, we additionally implement an ablation variant, Super-LIO*, which is identical to Super-LIO but replaces HKNN with the 18-neighbor voxel search of Faster-LIO. We also provide a concurrency-optimized version of FAST-LIO2, denoted FAST-LIO2*, which preserves the original algorithm but replaces the OpenMP-based parallel implementation in the official release with Intel TBB, ensuring that its concurrency behavior matches that of Super-LIO for fair runtime and resource-usage evaluation.
For fairness, all methods use identical parameters: maximum iterations = 4, random downsampling rate = 3, and voxel-filter resolution = 0.5 m. For voxel-based methods [faster_lio, ig_lio] including Super-LIO, the map voxel size is also set to 0.5 m. Super-LIO uses an HKNN radius of $R_{\text{max}}=0.875\,\mathrm{m}$ and searches within a $7\times 7\times 7$ subvoxel neighborhood. Except for dataset-specific LiDAR-IMU extrinsics, all parameters remain unchanged across experiments.
Experiments are conducted on both public and private datasets. Public benchmarks include M2DGR [m2dgr], NCLT [nclt], MCD [MCD], and NTU VIRAL [NTU], all providing ground-truth trajectories. Our private dataset contains ten isequences covering diverse indoor and outdoor environments, used for runtime and robustness tests. All data are replayed at $5\times$ speed on the X86 platform(AMD 5800H) to evaluate throughput and accuracy, and at $1\times$ speed on the NVIDIA Orin NX to evaluate real-time performance under resource constraints. Examples of private data collection platforms and sequences are shown in Fig. 4 and Fig. 5.
@@
Figure 5: Point cloud maps produced by Super-LIO on ten self-collected evaluation sequences. The sequences span diverse environments, including forests, parks, underground garages, and office buildings. “h” denotes data collected with a handheld device, and “u” denotes data collected on a UAV platform. These varied scenarios and motion conditions are used to evaluate computational efficiency and robustness.
IV-B Accuracy Evaluation
TABLE I: RMSE (m) on Public Datasets (X86 @ 5x Speed)